OpenStudy (anonymous):

In 1990 the life expectancy of males was 66.7 years. In 1995 it was 69.9 years. Let E be the life expectancy in year t and let t represent the number of years since 1990. What is the linear function E(t) that fits the data E(t)=?t + ? and using these functions predict the life expectancy of males in 2006. E(16) = ?

OpenStudy (anonymous):

you have two points on the line: (0,66.7) and (5,69.9). Find the slope, and use point slope form to give the equation of the line through these points. Then you can plug in t=16 to predict 2006.

OpenStudy (anonymous):

i am on the same problem with different numbers and I do not quiet under stand what you mean to put the point slope in the equation

OpenStudy (anonymous):

find the slope first: \[slope=\frac{\Delta y}{\Delta x} \] (69.9-66.7)/(5-0)=3.2/5 now we have the slope and a point. choose the point (0,66.7) eqn is: \[E(t)-66.7=\frac{3.2}{5}(t-0)\] \[E(t)=\frac{3.2}{5}t+66.7\] now you can plug in t=16